OpenStudy (anonymous):

x varies jointly as y and z. 1 set of values is x=42,y=3,z=3

OpenStudy (anonymous):

ok so this uses the equation x=kyz plug in the values to solve for k 42=k*3*3 k=42/9 so x=(42/9)y*z

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

i have alot more ? if u want to help me with them i promise to give u a medal for everyone u answer

OpenStudy (anonymous):

sure. Ill be here for a while.

OpenStudy (anonymous):

k thanks im typing one right now